Circular No. 8245

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003kf, 2003kg, 2003kh, 2003ki Four additional apparent supernovae have been found on unfiltered

CCD images — SN 2003kf by W. Li (LOSS/KAIT; cf. IAUC 8243), SNe 2003kg and 2003kh by M. Armstrong (cf. IAUC 8241), and SN 2003ki by T. Boles (cf. IAUC 8233).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003kf Nov. 27.4  6^{h}04^{m}35.^{s}42 −12^{o}37′42′′.8 14.8 9′′.2 E, 14′′.3 S
2003kg Nov. 27.79 22 03 34.64 +15 06 10.6 16.6 5′′.6 W, 15′′.5 S
2003kh Nov. 27.80 19 52 17.93 +58 58 39.9 17.5 10′′.9 W, 12′′.6 N
2003ki Nov. 28.00 7 51 33.24 +63 55 51.6 18.0 5′′.9 W, 9′′.6 N

Additional approximate magnitudes (by the respective discoverers unless otherwise noted): SN 2003kf in MCG –02-16-2, Nov. 18.4 UT, [19.0 (Li); 28.022, 14.3 (Armstrong; position end figures 35^{s}.47, 42′′.5); 28.540, 15.7 (K. Itagaki, Teppo-cho, Yamagata, Japan; 0.60-m reflector; position end figures 35^{s}.46, 42′′.5; communicated via S. Nakano, Sumoto, Japan). SN 2003kg in IC 1427, Oct. 16, [19.0; Nov. 5, [19.0; 28.723, 16.6. SN 2003kh in UGC 11491, July 19, [19.0; Nov. 12, [19.0; 28.720, 17.5. SN 2003ki IN MCG +11-10-34, Oct. 17, [19.5; 27, [19.5; Nov. 28.730, 18.0. Nothing was visible at the positions of SNe 2003kg and 2003kh on Palomar Sky Survey red and blue plates. SN 2003ki is not present on Digitized Sky Survey red (1998 Nov. 24) or blue (1997 Feb. 4) plates.

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 320–880 nm) of SN 2003kf, obtained by Matheson on Nov. 28.29 UT with the Baade 6.5-m Magellan telescope (+ IMACS spectrograph), shows it to be a type-Ia supernova near maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 2215 km/s for the host galaxy, the supernova expansion velocity is 12600 km/s for Si II (rest 635.5 nm).

SUPERNOVA 2003kd IN UGC 2468 Matheson et al. also communicate that a spectrogram (range 370–

750 nm) of SN 2003kd (cf. IAUC 8243), obtained by P. Berlind on Nov. 26.40 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova near maximum.

2003 November 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green